### v1

# MemoFurther Readings

Abstract
Coordinated information operations remain a persistent challenge on social media, despite platform efforts to curb them. While previous research has primarily focused on identifying these operations within individual platforms, this study shows that coordination frequently transcends platform boundaries. Leveraging newly collected data of online conversations related to the 2024 U.S. Election across $\mathbb{X}$ (formerly, Twitter), Facebook, and Telegram, we construct similarity networks to detect coordinated communities exhibiting suspicious sharing behaviors within and across platforms. Proposing an advanced coordination detection model, we reveal evidence of potential foreign interference, with Russian-affiliated media being systematically promoted across Telegram and $\mathbb{X}$. Our analysis also uncovers substantial intra- and cross-platform coordinated inauthentic activity, driving the spread of highly partisan, low-credibility, and conspiratorial content. These findings highlight the urgent need for regulatory measures that extend beyond individual platforms to effectively address the growing challenge of cross-platform coordinated influence campaigns.
Introduction
Social media has evolved into a vital arena for public discourse, serving as a platform where individuals and communities can converge to discuss political, social, and cultural issues. These platforms have been instrumental in fostering large-scale movements, such as the Arab Spring and Black Lives Matter, where activists and citizens alike have used social media to amplify calls for justice and change (Howard and Hussain, 2013; Freelon et al., 2016). However, social media has also been linked to polarization and radicalization, as exemplified by the events surrounding the U.S. Capitol attack (Election Integrity Partnership, 2020; Luceri et al., 2021), where online platforms played a role in mobilizing and coordinating participants  (Vishnuprasad et al., 2024). Similar cases include the Christchurch mosque shootings (Fakhrurroja et al., 2019) and far-right rallies across Europe (Klein and Muis, 2019; Pérez-Curiel, 2020), where social media content has been shown to exacerbate divisive sentiments and mobilize fringe communities (Davey and Ebner, 2019). The potential for influence of online discourse has contributed to the establishment of state-sponsored information operations aimed at steering public opinion by introducing false or misleading narratives on popular platforms (Nisbet and Kamenchuk, 2019; Wilson et al., 2018; Molter and DiResta, 2020).
Coordinated inauthentic behavior represents a prominent tactic for distorting public discourse by amplifying specific viewpoints and giving the illusion of widespread support (Giglietto et al., 2020; Khaund et al., 2021; Pacheco et al., 2021; Murero, 2023; Cima et al., 2024). These activities typically involve actions such as synchronized posting and co-activity behavior patterns (Oates and Gray, 2019; Pacheco et al., 2021; Luceri et al., 2024a)— such as similar retweets, hashtag sequences, link sharing— designed to push particular narratives to the forefront of public attention. Studies have also documented how these coordinated behaviors contribute to a range of detrimental social effects, such as the spread of propaganda (Hristakieva et al., 2022a), conspiracy theories (Tardelli et al., 2024), and the promotion of disinformation (Weber and Neumann, 2021). Moreover, coordinated online actions have been linked to heightened toxicity in online conversations (Loru et al., 2024) and the dissemination of extremist ideologies (Kursuncu et al., 2021; Vishnuprasad et al., 2024).
In recent years, a growing body of research has focused on the role of social media in the context of political elections. For example, studies on the Brazilian elections have documented the strategic use of bots and coordinated networks to shape public opinion and influence voter perceptions, highlighting the risks of computational propaganda in electoral processes (Pacheco, 2024). Similarly, research on European elections has illustrated how misinformation campaigns spread across social media can alter public perceptions, as seen in countries such as France, Germany, and Italy (Ferrara, 2017; Lovari, 2020; Baqir et al., 2024; Nogara et al., 2024). In the United States, coordinated online activities around elections have garnered significant attention, with investigations revealing how coordinated bot activities and foreign information operations have sought to manipulate voter beliefs and amplify divisive content across multiple election cycles (Bessi and Ferrara, 2016; Howard et al., 2018; Luceri et al., 2019b; Tardelli et al., 2024). Notably, this large body of literature has predominantly focused on identifying such operations within individual platforms.
Contribution of this work
In this work, we expand on current research in coordination detection by broadening our scope to include the identification of coordinated inauthentic activity (CoIA) spanning multiple platforms. We build on, adapt, and advance state-of-the-art coordination detection techniques to identify intra- and cross-platform CoIA that promotes external web domains and amplify specific narratives in the context of the 2024 U.S. Presidential Election. We leverage a large-scale dataset covering election-related online conversations spanning several platforms, including Twitter/$\mathbb{X}$, Facebook, and Telegram, collected during the run-up months to the Presidential Election. Combining insights from computational techniques for CoIA detection and language models for content analysis, we aim to answer the following Research Questions (RQs):
  • RQ1:
    Web Domain Promotion. Do we observe intra- and cross-platform CoIAs aimed at redirecting traffic towards specific web domains? What are the characteristics of coordinated accounts, and which specific domains do they promote?
  • RQ2:
    Content Amplification. Do we observe CoIAs pushing specific narratives? What are the characteristics of coordinated accounts, and which specific topics do they amplify?
  • RQ3:
    Engagement & Impact. What level of engagement do CoIAs generate across various platforms? And how does it compare to the engagement garnered by organic users?
Leveraging our multi-platform dataset and advancing CoIA detection techniques, we uncover multiple networks of coordinated inauthentic accounts. We analyzed the textual content and link-sharing behaviors of these accounts, identifying the narratives they aim to amplify and the external domains they direct traffic to, assessing both their credibility and ties to foreign and domestic entities. Our findings reveal coordinated efforts to promote Russian-affiliated media across Telegram and $\mathbb{X}$, with highly partisan, low-credibility content systematically amplified by these networks. Conspiracy theories surrounding public health, the environment, and political topics, such as immigration and geopolitical tensions, were prominently featured. Notably, QAnon-related narratives were especially prevalent on Telegram, with coordinated accounts driving much of the discussion. Furthermore, we assessed the prevalence of AI-generated content produced by coordinated actors and the level of engagement their content attracted. We found that coordinated actors on Telegram relied on AI-generated content significantly more than organic users, while the opposite trend was observed on Facebook. This study sheds new light on cross-platform coordination efforts related to the upcoming U.S. Presidential Election, revealing the complex dynamics of influence operations. These findings underscore the urgent need for regulatory measures that go beyond individual platforms to effectively tackle the growing challenge of cross-platform coordinated inauthentic activity.
Related Work
Influence Campaigns Interfering with U.S. Elections
A growing body of research has focused on the influence of coordinated disinformation campaigns during U.S. elections, particularly through the manipulation of social media platforms (Sharma et al., 2022; Ferrara et al., 2020). Studies have uncovered significant efforts by foreign and domestic actors to manipulate the political landscape during both the 2016 and 2020 U.S. Presidential Elections (Bessi and Ferrara, 2016; Vishnuprasad et al., 2024). For example, the Russian Internet Research Agency (IRA) was linked to a large-scale disinformation campaign during the 2016 election, deploying thousands of social bots and state-sponsored trolls to promote divisive narratives and foster discord (Badawy et al., 2019, 2018). The activity of these inauthentic actors has been extensively studied on Twitter (Zannettou et al., 2020; Luceri et al., 2020; Addawood et al., 2019; Luceri et al., 2019a), with fewer studies examining their presence on other platforms (Zannettou et al., 2019b, a).
Similarly, the 2020 U.S. Presidential Election witnessed the propagation of various false claims and conspiracy theories, including narratives about voter fraud and COVID-19 misinformation (Ferrara et al., 2020). Prominent examples include the “Stop the Steal” movement, which spread across platforms like Twitter and Facebook, inciting allegations that the election results were fraudulent (Election Integrity Partnership, 2020). Coordinated activity surrounding this narrative was widespread, often involving amplification techniques such as automated retweets or sharing similar content across accounts to create an illusion of widespread consensus (Luceri et al., 2021).
Detection of Coordinated Inauthentic Activity
Coordinated inauthentic activity has been identified as a predominant tactic in spreading disinformation and conspiratorial content (Luceri et al., 2024a; Pacheco et al., 2020). Researchers have developed sophisticated methods to detect coordination. Coordination is not limited to malicious campaigns: it can also encompass legitimate organizing efforts, social movements, etc. However, in the context of disinformation campaigns, coordinated behavior is typically characterized by manipulative actions aimed at amplifying false narratives (Pacheco et al., 2021; Nizzoli et al., 2021).
The detection of such deceptive, orchestrated efforts has evolved to address both automated and human-coordinated activities through advanced machine learning and network-based approaches. Machine learning techniques have traditionally focused on identifying automated activities, such as bot-like behavior, and distinguishing them from human actions (Yang et al., 2019; Cresci et al., 2016). Recently, attention has shifted toward characterizing suspicious human-operated accounts, with research emphasizing content-, behavioral-, and sequence-based methods to detect coordinated actions by state-sponsored trolls (Hristakieva et al., 2022b; Alizadeh et al., 2020; Luceri et al., 2024, 2020; Nwala et al., 2023; Ezzeddine et al., 2023).
Network-based detection, however, has gained prominence due to its ability to reveal coordinated behavior by constructing networks that highlight similarities in user actions, such as shared content, hashtags, or synchronized posting times (Pacheco et al., 2020; Pacheco et al., 2021; Nizzoli et al., 2021; Mannocci et al., 2024; Magelinski et al., 2022; Luceri et al., 2024a). These networks are analyzed using properties like node centrality and edge weight, which help identify clusters of users involved in coordinated IOs (Luceri et al., 2024a; Vishnuprasad et al., 2024; Tardelli et al., 2023). This approach has proven effective in uncovering both automated and human-coordinated activities, providing critical insights into the tactics and structure of influence campaigns.
In contrast to previous cross-platform studies (Ng et al., 2022; Gatta et al., 2023; Zannettou et al., 2019a), our approach introduces a more advanced and fine-grained coordination detection framework, specifically crafted to minimize false positives by combining network measures (edge weights and node centralities) to identify a highly coordinated set of inauthentic accounts across multiple platforms. Unlike existing methods, which have explored dynamics between mainstream and fringe platforms, our work leverages state-of-the-art language models and AI-generated content detectors to provide a nuanced characterization of users across a broader selection of mainstream social media platforms.
Data
Collection. The rationale behind our data collection is to capture the online discourse surrounding the 2024 U.S. Presidential Election across multiple social media networks. The data collection for our study spans May and June 2024 and covers three major online platforms: Facebook, $\mathbb{X}$ (formerly Twitter), and Telegram. Each platform was queried to obtain posts or messages containing specific election-related keywords. The full list of keywords used to filter and collect data is provided in Appendix 1.
Facebook data was collected through (the now defunct) Crowdtangle, which offered access to public posts of groups and pages. For $\mathbb{X}$, we gather publicly available information, including original tweets, retweets, replies, and quotes, retrieved via the platform’s web interface. Finally, Telegram data was gathered using the Telegram API, allowing the extraction of public chats’ details, meta data, messages, and message attachments. Details on the $\mathbb{X}$ and Telegram data collection infrastructure can be found in (Balasubramanian et al., 2024a; Blas et al., 2024b).
Platform Accounts Posts/Tweets/Messages URLs Domains
Facebook 6,137 46,310 15,009 5,247
$\mathbb{X}$/Twitter 178,379 6,021,428 582,052 35,922
Telegram 15,537 4,309,880 2,087,078 183,924
Table 1. Datasets statistics.
Statistics. For each platform, the collected content consists of posts on Facebook, tweets on $\mathbb{X}$, and messages on Telegram. To identify potential coordination among accounts operating within each platform, we analyze the textual similarity of their shared content as well as similarity patterns in co-URL sharing. To establish coordination across platforms, we examine only the co-sharing of URLs, as they represent specific pieces of information that can be easily tracked across different platforms (Luceri et al., 2024b). In fact, URLs can be embedded in user’s posts on each platform under analysis. These URLs serve as a common thread linking sharing activities across the Web. Thereby, we identify and count unique URLs embedded within the shared content of each dataset. Table 1 presents a summary of the datasets, including the number of accounts, [^1] We will use the term accounts to also refer to Facebook pages, Facebook groups, and Telegram channels. posts, and unique URLs for each platform.
Methodology
Exposing CoIA within and across Social Media Platforms
Coordinated accounts can execute their campaigns using various strategies; here, we analyze campaigns that focus on promoting specific URLs (or web domains) and campaigns that produce highly similar textual content to amplify certain topics. The former strategy is commonly used to create the illusion of public consensus around certain viewpoints by artificially amplifying links to external webpages, mock websites, and other social media networks (Minici et al., 2024a; Gabriel et al., 2023). The latter is often employed to manipulate platform feed algorithms by pushing specific keywords or hashtags, attempting to boost trending topics, and making content appear more popular than it actually is (Pacheco et al., 2021; Vishnuprasad et al., 2024). We refer to these orchestrated efforts as web domain promotion and content amplification. Since URL-sharing functions through a consistent mechanism across all observed platforms, we analyze web domain promotion both within and across platforms. Due to substantial variability in textual content-sharing—driven by platform-specific constraints, such as character limits on some platforms, and differences in the nature of messaging (e.g., Telegram chats versus Facebook and $\mathbb{X}$ posts)—our analysis of content amplification is focused on within-platform activity.
Detection of Web Domain Promotion
To identify orchestrated campaigns promoting specific web domains, we examine user similarities based on the URLs they share in their social media posts. To ensure high-quality data and reduce noise, we applied several preprocessing steps. First, we imposed a minimum activity threshold, requiring each user to have shared at least 10 unique URLs. This criterion is consistent with prior work (Luceri et al., 2024a; Pacheco et al., 2021) and ensures that the analysis focuses on users with substantial contributions, filtering out those with minimal engagement that could distort the detection of coordinated accounts. Next, we expanded shortened or obfuscated URLs using a URL expansion library [^2] https://github.com/dfreelon/unspooler.
Co-URL similarity network. We constructed a bipartite user-URL network, where users and URLs are connected based on the URLs that users share. In the adjacency matrix of the bipartite graph, users are represented by the rows, whereas the columns represent the URLs. In accordance with previous work (Luceri et al., 2024a; Pacheco et al., 2021), we applied the Term Frequency-Inverse Document Frequency (TF-IDF) transformation to represent the user-URL matrix. To avoid bias towards overly frequent URLs, we set a maximum document frequency (max DF) of the 90th percentile and a minimum document frequency (min DF) threshold of 5 occurrences per URL. This ensures that both very rare and overly common URLs are excluded from the analysis, leaving us with a meaningful set of URLs that can identify CoIAs.
The similarity between users is built by comparing their pairwise co-shared URL patterns. This is obtained by projecting the bipartite network into a user-to-user similarity network. For network construction, we computed exhaustive pairwise cosine similarity between user vectors of the bipartite graph. This similarity measure captures how similar users are based on the URLs they shared, with higher similarity scores indicating stronger coordination. This co-URL similarity network forms the foundation to identify coordinated users potentially driving campaigns aimed at redirecting traffic to specific web domains.
Intra-platform and cross-platform detection. Being agnostic to the platform under scrutiny, URLs serve as common entities that can be analyzed across different platforms to identify potential cross-platform campaigns. Therefore, we constructed two types of co-URL networks: intra-platform and cross-platform. For the intra-platform network, we computed cosine similarity between all pairs of users within the same platform. For the cross-platform network, we calculated pairwise similarity only between users from different platforms, linking users based on their shared URL patterns.
Extraction of network properties for coordinated accounts detection. To identify online coordination, we build upon and integrate state-of-the-art strategies that filter either low-weight edges or peripheral nodes from the similarity network to detect CoIAs. The first approach (Pacheco et al., 2021) prioritizes the strength of similarity, filtering out low-weight edges likely representing spurious similarities. The second approach (Luceri et al., 2024a) focuses on the breadth of similarities, pruning nodes based on their centrality. Luceri et al. (Luceri et al., 2024a) demonstrated that eigenvector centrality is an effective network property for identifying the most suspicious users in a similarity network, where coordinated accounts tend to share numerous similarities with other highly connected nodes. The assumption is that coordinated activity typically involves multiple accounts. In a similarity network, coordinated actions manifest as a pronounced collective similarity, where a coordinated account is highly connected (i.e., similar) to numerous other nodes that are themselves well-connected. Consequently, eigenvector centrality has proven to be an effective network property for detecting online coordination (Luceri et al., 2024a). Here, we extend this notion of collective similarity by considering the density of the similarity network. Network density is a measure that quantifies how close a network is to being fully connected and, in this scenario, it provides an exact measure of collective similarity.
In particular, we use network density as a variable to modulate the combination of thresholds for both node and edge filtering. By integrating these strategies, we aim to leverage the strengths of both filtering techniques, constructing a model that accounts for both the breadth and strength of similarities. This integrated approach is particularly important, as these techniques have not yet been applied to platforms like Telegram and Facebook, where coordination dynamics may vary significantly.
Density-based unsupervised network dismantling. To combine these strategies in a fully unsupervised manner, we developed a novel method for conducting a grid search of parameters in the two-dimensional space defined by node centrality and edge weight distributions. Specifically, the grid search explores two parameters—the quantile of node centrality and the quantile of edge similarity—and evaluates network density for each combination and every connected component in the similarity graph. We hypothesize that, as we progressively filter the similarity graph by removing edges and nodes, a transitional phase in the density of the connected components will signal potentially coordinated accounts. To ensure robustness, we adopt a conservative approach by evaluating the transitional phase of the smallest density among all connected components in the filtered similarity graph. By controlling the component with the lowest density, we ensure that all other components exhibit higher densities. This conservative choice is driven by the unsupervised nature of the task, prioritizing the minimization of false positives, i.e., legitimate users misclassified as coordinated. Therefore, filtering thresholds were chosen based on the transitional phase of the smallest density of a connected component of the similarity graph, thus, focusing on identifying highly suspicious CoIA. Details of the parameters used for detection are provided in the Appendix.
Detection of Content Amplification
Coordinated actors may employ a variety of tactics to achieve their goals. A common tactic is to artificially amplify content on specific topics to create the appearance of widespread grassroots support and manipulate platforms’ feed algorithms (Mannocci et al., 2024; Vishnuprasad et al., 2024; Pacheco et al., 2020). To uncover content amplification, we construct a Text Similarity Network (TSN), where nodes represent users linked by the similarity of the content they share. This TSN is then employed to identify a subset of users exhibiting suspiciously coordinated behavior.
The initial step in constructing the TSN involves preprocessing the raw data to ensure the results are meaningful. In line with standard practices (Luceri et al., 2024a), we exclude retweets and remove punctuation, stopwords, emojis, URLs, as well as any content with fewer than four words. To capture the semantic nuances of the content, we embed all text data using the SentenceTransformer model stsb-xlm-r-multilingual [^3] https://huggingface.co/sentence-transformers/stsb-xlm-r-multilingual, and then calculate the average cosine similarity between pairs of users. To avoid spurious correlations, we consider text similarity valid only if it occurs within a one-day sliding window. For each window, we calculate the text similarity between the shared content of user pairs, setting the edge weight in the TSN to the average similarity observed across all windows.
To identify coordinated accounts, we apply established thresholds from the literature (Pacheco et al., 2021; Luceri et al., 2024a). Specifically, we filter out edges with similarity below 0.95 and classify users as coordinated if their nodes rank in the top 0.5% by eigenvector centrality.
Characterizing Content Pushed by CoIA
In this section, we present an overview of the methods used to characterize textual content amplified by coordinated actors across various social media networks. This characterization encompasses topic analysis, AI-generated content detection, and an assessment of content credibility.
Topic Analysis
Coordinated accounts often amplify specific narratives or themes to steer public attention toward polarized, inflammatory, or misleading discussions. To uncover the agendas these accounts seek to promote, we use BERTopic (Grootendorst, 2022), a state-of-the-art tool for topic extraction. For a detailed description of BERTopic’s methodology, we refer readers to (Grootendorst, 2022). This approach also helps identify patterns in shared content, improving our understanding of coordinated activities and their intended impact on public discourse.
Once coordinated accounts are identified using the methodology outlined in Sec. Detection of Content Amplification, we apply BERTopic to the entire content corpus. This allows us to map the topics promoted by coordinated accounts and compare them with those shared by organic users (i.e., users not classified as coordinated). We choose BERTopic over alternatives such as LDA or GPT-based approaches because it offers an effective balance between accuracy and scalability.
AI-Generated Content Detection
To identify AI-generated text, we use the approach introduced by (Dmonte et al., 2024), which is designed to identify AI-generated content on $\mathbb{X}$. We initially trained a RoBERTa [^4] https://huggingface.co/FacebookAI/roberta-base model on their $\mathbb{X}$ dataset and then evaluated its performance on datasets from other platforms by developing a new validation set containing approximately 2,000 samples. This validation set was constructed using Llama 3.1 3B Instruct [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct, GPT-4o [^6] https://openai.com/index/hello-gpt-4o/, and open-source social media datasets from Telegram, $\mathbb{X}$, and Facebook, sampled from posts between 2010 and 2015. We assume these older datasets do not contain AI-generated content and sampled around 1,000 sentences from them. To create a balanced set of human and AI-generated content, we prompted the language models to generate 1,000 samples that mirrored the original texts in topic distribution and length. This approach ensured that the AI-generated texts were comparable to human-generated samples in both thematic and structural characteristics. Results showed precision values ranging from 0.87 to 0.97 in detecting AI-generated content within the validation set, aligning with our goal of prioritizing precision over recall to minimize false positives (misclassified legitimate users). Full results are available in the Appendix.
Credibility Assessment
We assess the credibility of web domains using Media Bias/Fact Check (MBFC). MBFC is an independent watchdog that rates news outlets on a 6-point factuality scale, ranging from Very Low to Very High. For each post in our datasets, we systematically extract, expand, and parse all embedded URLs, checking whether the URL belongs to a low- or high-credibility domain from our lists.
Given recent news [^7] https://www.bbc.com/news/articles/c8rx28v1vpro [^8] https://www.state.gov/u-s-department-of-state-takes-actions-to-counter-russian-influence-and-interference-in-u-s-elections/ and prior instances of interference by Russian agencies in U.S. elections (Badawy et al., 2019), we also examined the potential amplification of Russian media outlets. Following a similar approach to previous work (Pierri et al., 2023), we utilize the VoynaSlov dataset (Park et al., 2022) to obtain a list of 23 state-affiliated Russian websites. We then check whether the extracted URLs link to one of these web domains.
Finally, we employ a similar approach to identify content linked to conspiracies, such as QAnon, given the relevance of fringe theories during the 2020 US Election, and its aftermath (Sharma et al., 2022; Vishnuprasad et al., 2024). Building on the methodology of (Wang et al., 2023), we detect posts sharing QAnon content by utilizing a list of keywords commonly associated with the conspiracy, as outlined by (Sharma et al., 2022).
Results
This section presents the results obtained from the detection of online coordination. We first present the results for the detection of web domain promotion and content amplification. Following the identification of these campaigns, we examine the content pushed by coordinated networks and the engagement received by suspicious actors focusing on several interaction and engagement metrics.
Detecting CoIA for Web Domain Promotion (RQ1)
We detect CoIAs that push web domains within and across platforms as described in Section Detection of Web Domain Promotion. We begin by presenting results for each coordinated network within individual platforms. Following this, we analyze the inter-platform CoIA.
To provide a comprehensive view, we created a cross-platform network by combining the intra- and inter-platform coordinated networks into a unified network. In this network, nodes are connected if they are linked in either the intra- or inter-platform coordinated networks. The primary result of this analysis is displayed in Figure 1, illustrating coordinated activity both within and across the three platforms under study.
Figure 1. Cross-platform coordination network showing user coordination across four social media networks.
Intra-Platform Coordination.
We first analyze each intra-platform CoIA based on the co-URL similarity networks extracted from each platform separately. We describe intra-platform CoIAs as follows.
Telegram: We identified 33 highly coordinated channels co-sharing URLs to web domains with a partisan slant and low factuality. As shown in the Table 2, the most frequently shared domains are predominantly right-leaning and of low credibility, as assessed by MBFC [^9] https://mediabiasfactcheck.com/. Notably, the table highlights the presence of a Russian state-controlled media outlet (RT.com) and its video-on-demand subsidiary (Ruptly.tv), which could indicate a foreign effort to interfere in the election. Both RT.com and Ruptly.tv have previously been accused of orchestrating campaigns to influence U.S. elections via social media [^10] https://www.nbcnews.com/politics/2020-election/facebook-blocks-russia-backed-accounts-other-sites-keep-churning-out-n1242683.
Additionally, this CoIA promotes a far-right website (thegatewaypundit.com) and an extremist-friendly video platform (odysee.com), suggesting ties to fringe ideas. A manual review of these Telegram channels, including their profile descriptions (see Table 8) and shared messages (see Table 12), reveals a strong prevalence of content and accounts promoting conspiracy theories, particularly those related to COVID vaccines, 5G, and alternative news media.
Domain Shares Factuality Leaning
ruptly.tv 2,117 Mixed RIGHT-CENTER
rt.com 1,941 Very Low RIGHT-CENTER
odysee.com 1,602 Low RIGHT CONSPIRACY
dailymail.co.uk 1,159 Low RIGHT
thegatewaypundit.com 746 Very Low EXTREME RIGHT
Table 2. Top-5 domains shared in Telegram by coordinated accounts. Columns represent, from left to right: domain, the number of shares among coordinated accounts, and the factuality and political leaning scores based on data from MBFC.
$\mathbb{X}$/Twitter: We identified a network of 19 coordinated accounts predominantly promoting content from right-leaning domains, as listed in Table 3. A manual review of this Twitter CoIA reveals that many of these coordinated accounts share similar profile descriptions, reflecting narratives tied to religious and conservative principles (see Table 9). Additionally, an analysis of the messages (see Table 13) shows a strong presence of politically partisan content, predominantly aligned with right-leaning ideologies.
Domain Shares Factuality Leaning
foxnews.com 880 Mixed RIGHT
foxbusiness.com 13 Mixed RIGHT-CENTER
newsbreakapp.com 5 NA NA
go.shr.lc 2 NA NA
lifenews.com 2 Low FAR RIGHT
Table 3. Top-5 domains shared in Twitter by coordinated accounts. Columns represent, from left to right: domain, the number of shares among coordinated accounts, and the factuality and political leaning scores based on data from MBFC.
Facebook: On Facebook, no well-known media outlets dominate the shared content. However, the most shared domain is a partisan news outlet, as suggested by the semantics of its title: gorightnews.com. Additionally, there is a direct connection to a public activist known for political campaigns [^11] https://en.wikipedia.org/wiki/Peter_Boykin and their website: peterboykin.com. These campaigns and related Facebook accounts primarily focus on the “Gays for Trump" movement, an American LGBTQ organization that advocates for former U.S. President Donald Trump and his administration  [^12] https://en.wikipedia.org/wiki/Gays_for_Trump. The top-engagement messages (see Table 14) and bios (see Table 10) indicate strong support for right-leaning ideologies.
Cross-Platform Coordination.
The analysis of the cross-platform network reveals distinct patterns. The unified cross-platform network, displayed in Fig. 1, highlights a giant component connecting coordinated accounts from both Telegram and Twitter, while the Facebook coordination network remains largely disconnected. A few noteworthy observations:
First, we observe that most cross-platform connections occur between Twitter accounts and Telegram channels, likely due to their higher representation in the dataset. Examining the bios of the top users by degree, as shown in the Appendix (see Table 15), we note a prominent presence of non-mainstream news outlets and conspiracy theories, such as the flat earth one.
Second, when focusing on the bridge nodes between the Telegram-coordinated cluster and the cross-platform giant component, we find highly influential channels with up to 24,000 subscribers. These channels often feature bios referencing free speech and religion, such as:
  • "The FIGHT for FREEDOM - :Aron TRUTH Social"
  • "Q Reee-searchers Watchers"
  • "’Arise and shine, for your light has come, and the glory of the LORD rises upon you.’ Isa. 60:1 WE RISE beyond the challenges of yesterday to become better stewards of tomorrow!"
Similarly, the bridge node between the Twitter-coordinated cluster and the cross-platform giant component is a highly influential account with 32,000 followers. The bio reads: "NO DM’S. Beautiful disaster. Self-proclaimed arbiter of great ideas. Here to annoy the dumb asses. NO LISTS!! #Imvotingforafelon #animallover", indicating a strong partisan flavor.
Finally, the largest component of the cross-platform coordination network is dominated by a mix of Telegram and Twitter accounts promoting domains such as magapac.com, QAnon-related narratives (e.g., "WWG1WGA" [^13] https://en.wiktionary.org/wiki/WWG1WGA), and accounts with partisan bios like:
  • "We The Ultra Patriots is made up of patriotic Americans dedicated to exposing crimes against humanity, false flags, lies, and corruption. We The Ultra Patriots are…"
A closer look at the top shared domains within the coordinated network (see Table 4) reveals that most are partisan, including truthsocial.com, an alternative non-mainstream social media platform [^14] https://en.wikipedia.org/wiki/Truth_Social. Additionally, seven of the shared domains are Russian state-affiliated websites [^15] https://github.com/chan0park/VoynaSlov/tree/master, including tv5, rbc, gazeta, ruptly, and mil, further suggesting potential Russian information operations to interfere with the election discourse. Examining the posts and messages with the highest engagement (see Table 15 in the Appendix), we observe the prevalence of ultra-MAGA narratives, conservative religious themes, and environmental news.
Domain Shares Factuality Leaning
thegatewaypundit.com 23,513 Very Low EXTREME RIGHT
zerohedge.com 8,331 Low RIGHT CONSPIRACY
truthsocial.com 7,800 NA NA
nypost.com 6,797 Mixed RIGHT-CENTER
theepochtimes.com 6,241 Mixed RIGHT
Table 4. Top domains shared by cross-platform coordinated accounts. Columns represent, from left to right: domain, the number of shares among coordinated accounts, and the factuality and political leaning scores based on data from MBFC.
(a) Telegram
(b) Twitter
(c) Facebook
Figure 2. Content analysis of coordinated channels on Telegram, Twitter, and Facebook. We report the top-10 most recurring topics and compare their prevalence against the organic discourse.
Detecting CoIA for Content Amplification (RQ2)
In this section, we present our findings on the specific topics amplified by coordinated users within content amplification campaigns. Following the method in Sec. Detection of Content Amplification, we extract the Text Similarity Network and identify a seed set of coordinated users. We then apply BERTopic, as described in Sec. Topic Analysis, to identify the themes these users aim to promote and compare them to the themes that emerge organically on each platform. For each platform, we report the top-10 most prevalent topics within the coordinated cohort and compare their prevalence to that within the organic population.
This approach highlights themes promoted by coordinated users, with large differences indicating topics characteristic of coordinated activity. In the remainder of this section, we detail our findings on coordination-driven content amplification across the three platforms analyzed.
Telegram: Based on the significantly high similarity of their shared content, we identified 57 coordinated Telegram channels. Among these channels, public health is a prominent topic of discussion, as can be seen in Figure 2(a), with skepticism around the COVID-19 vaccine emerging as the most frequently discussed theme. Additionally, many messages promote alternative medicine, and there is notable discussion surrounding a conspiracy theory related to bird flu vaccination. Telegram’s coordinated channels also focus on strictly political discussions, particularly on the U.S. presidential debate, the Russia-Ukraine conflict, and Donald Trump’s legal challenges. Interestingly, we observe discourse around immigration issues in Ireland, which appears to be echoed by accounts associated with the MAGA movement. Notably, three out of the ten topics involve conspiracy theories related to flat-earth beliefs, bird flu, and antisemitism. The most representative content for each topic, as identified by BERTopic, is provided in the Appendix.
We also assess the presence of QAnon-related keywords in content shared by both coordinated and organic channels. Approximately 1.66% of content from coordinated users contains at least one QAnon-related keyword, compared to only 0.12% of content from the organic population. This substantial difference is evident in Table 5, where we list the top five keywords present in the coordinated group, also compared to keywords shared by organic channels. Although coordinated channels make up only 0.36% of the total Telegram dataset, they show a marked prominence, both in absolute and relative terms, in sharing QAnon-related keywords compared to the organic channels. These findings indicate a propensity for Telegram CoIAs to disseminate conspiracies.
Keyword Count / Prevalence
Coordinated Organic
wwg1wga 1260 (0.59%) 520 (0.05%)
plandemic 619 (0.29%) 102 (0.01%)
adrenochrome 448 (0.21%) 118 (0.01%)
qanon 260 (0.12%) 107 (0.01%)
deepstate 193 (0.09%) 87 (0.008%)
Table 5. Keyword statistics in coordinated and organic content related to QAnon conspiracies on Telegram.
$\mathbb{X}$/Twitter: On Twitter, we identified 221 coordinated users. Figure 2(b) displays the top 10 topics shared by these coordinated accounts, all of which pertain to social, economic, or political issues. The most prevalent topic — and one that diverges significantly from those shared by the organic group — focuses on the Independent candidate Robert F. Kennedy, with a subset of coordinated accounts promoting him as a serious contender in the U.S. Presidential Election. Coordinated users also concentrate on the Democratic Primary race in Westchester County (labeled as “bowman-lattimer” in the figure), where the most representative tweets endorse candidate Lattimer. This discussion appears to be framed around the anti-Israel stance of the incumbent, Bowman. The remaining topics revolve around polarizing issues such as COVID-19 vaccines, U.S. border security, and various scandals, including the Epstein files and the second son of Joe Biden (Hunter Biden). Also, these users discuss perceived opposition within the Black community to President Biden, critiques of the student loan forgiveness policy, and concerns over the Biden administration’s sale of oil reserves. No conspiracy theories were detected among the coordinated users, and only a minimal amount of QAnon-related content was found within the organic group. The most frequent keyword, “qanon”, appeared just five times.
Facebook: We identify 16 coordinated users in the Facebook discourse. As shown in Figure 2(c), most topics promoted by these users are left-leaning. A manual inspection reveals that all 16 accounts are associated with The Young Turks network [^16] https://en.wikipedia.org/wiki/The_Young_Turks, which MBFC classifies as left-leaning [^17] https://mediabiasfactcheck.com/the-young-turks/. Apart from a news item quoting a GOP representative accusing Biden of using “supplements” (i.e., “jacked-up Biden”), we found no trace of misleading, conspiratorial, or inflammatory content.
(a) Telegram
(b) Twitter
(c) Facebook
Figure 3. Empirical Cumulative Distribution of the generated engagement by coordinated and organic users.
Generated Engagement and AI-Generated Content
To comprehensively assess the impact of coordination strategies in the 2024 U.S. election online debate, we analyze the level of engagement generated by coordinated actors and the extent of their reliance on AI-generated content (AIGC).
Figure 4. Distribution of AIGC shared by coordinated and organic users across the three different platforms.
For each platform, we begin by merging the sets of coordinated actors identified through co-URL sharing (web domain promotion) and content similarity patterns (content amplification). We observe that the intersection between these two patterns is modest, with overlap rates of 1.3%, 0.5%, and 0% for Telegram, Twitter, and Facebook, respectively. This is in line with previous findings showing that different strategies are used to push diverse agendas through distinct coordinated networks (Ng and Carley, 2022; Luceri et al., 2024a). Aggregating coordinated actors from distinct CoIA networks, we obtain 233 coordinated actors on Telegram, 764 on Twitter, and 25 on Facebook.
Figure 3 shows the distribution of engagement generated by coordinated and organic users across all platforms. In each case, coordinated users tend to generate less engagement than their organic counterparts. This is expected, as the organic group includes major media outlets and influencers that significantly boost overall engagement. However, it is noteworthy that the engagement from coordinated actors is substantial, and particularly on Twitter and Facebook, the engagement distributions of coordinated and organic users are quite similar.
We also assess the prevalence of AIGC that was shared across the three platforms by both coordinated and organic users. These results are shown in Figure 4. Facebook is the social media platform where most AIGC is produced by both coordinated and organic accounts. While the AIGC prevalence is limited, we observe a striking difference in the activity between coordinated and organic users on Telegram and Facebook, whereas it is quite balanced on Twitter. Interestingly, AI-generated content is predominantly diffused by coordinated actors on Telegram, while an opposite trend is observed on Facebook.
Conclusions
This paper introduces a novel, network-based framework for detecting coordinated inauthentic activity (CoIA) across multiple social media platforms. Unlike most traditional methods that focus on isolated platform-specific analyses, our approach emphasizes cross-platform behaviors by constructing similarity networks that capture community-wide patterns of content sharing. By prioritizing both intra- and cross-platform connections, we are able to detect subtle yet significant coordination signals that remain hidden when platforms are analyzed in isolation.
Our unsupervised, network-based methodology allows us to identify a range of influence campaigns aimed at directing traffic toward specific narratives and domains. These campaigns frequently promote content that is partisan, low-credibility, and conspiratorial in nature. The model not only detects domestic actors but also reveals the systematic promotion of foreign-affiliated media, particularly Russian state-sponsored outlets, across platforms like Telegram and $\mathbb{X}$. This evidence points to a coordinated effort to amplify certain messages and domains across distinct platforms, particularly in the context of the 2024 U.S. Election, underscoring the necessity of a unified, cross-platform approach to tackle such influence operations (Gatta et al., 2023).
Limitations
While our study demonstrates substantial strengths, a few considerations remain. First, although the model leverages data from Telegram, Facebook, and $\mathbb{X}$, the specific timeframe under observation and platform-specific mechanisms may influence the broader applicability of our model and the generalizability of our findings to other scenarios and social media networks. Second, combining similarity networks with differing similarity distributions across platforms may introduce biases in the overall detection of coordination, potentially favoring certain platforms. Finally, this work does not examine multimedia content, thereby overlooking coordination across different data modalities and the potential use of generative AI techniques for producing synthetic media.
Future work should address these limitations by incorporating sophisticated statistical models to assess the significance of coordination patterns, thereby providing a more robust and statistically sound methodology to ensure that identified coordination signals are not merely artifacts of underlying data distributions. Expanding the framework to encompass additional platforms, data modalities, and a broader range of behavioral signals would further strengthen its capability to detect diverse forms of coordinated inauthentic activity.
About the Team
[…]
Appendix
1. Dataset Collection
[Uncaptioned image] [Uncaptioned image]
(a) Facebook posts (b) Facebook characters
[Uncaptioned image] [Uncaptioned image]
(c) Telegram posts (d) Twitter characters
[Uncaptioned image] [Uncaptioned image]
(e) Twitter posts (f) Telegram characters
Table 6. Distributions of posts and characters across platforms.
Keywords. Joe Biden, Donald Trump, 2024 US Elections, US Elections, 2024 Elections, 2024 Presidential Elections, Biden, Joe Biden, Joseph Biden, Biden2024, Donald Trump, Trump2024, trumpsupporters, trumptrain, republicansoftiktok, conservative, MAGA, KAG, GOP, CPAC, Nikki Haley, Ron DeSantis , RNC, democratsoftiktok, thedemocrats, DNC, Kamala Harris, Marianne Williamson, Dean Phillips, williamson2024, phillips2024, Democratic party, Republican party, Third Party, Green Party, Independent Party, No Labels, RFK Jr, Roberty F. Kennedy Jr. , Jill Stein, Cornel West, ultramaga, voteblue2024, letsgobrandon, bidenharris2024, makeamericagreatagain, Vivek Ramaswamy, JD Vance, Assassination, Tim Walz, WWG1WGA.
Additional dataset statistics. We present the distribution of the number of posts per user and post’s length for each platform.
2. AI Detection Classifier Validation
We utilized the AI-generated text detection classifier from (Dmonte et al., 2024), originally trained to identify AI-generated content in tweets. To extend the applicability of this classifier to other platforms, we constructed an external validation set using diverse, older datasets. Specifically, we built upon four datasets: Twitter-2010181818https://archive.org/details/twitter_cikm_2010, Facebook [^19] https://www.kaggle.com/datasets/sheenabatra/facebook-data, and Telegram.
From each of these datasets, we sampled 500 texts ranging in length from 125 to 1000 characters. Given the release years of these datasets, we assume that all texts are non-AI generated. To create the AI-generated class, we used GPT-4o and Llama 3.1 3B Instruct, prompting them to generate new texts with similar topics and lengths as the original non-AI texts. We performed this generation for each non-AI text and sampled 250 AI-generated texts from GPT and 250 from Llama.
Thus, for each platform, the validation dataset consists of 500 non-AI generated texts paired with 500 AI-generated texts: 250 from GPT and 250 from Llama, with both AI-generated sets matching the original texts in topic and length.
To further enhance the generalizability of the validation dataset, we incorporated non-AI generated content from additional sources: Reddit comments [^20] https://zissou.infosci.cornell.edu/convokit/datasets/reddit-coarse-discourse-corpus/, IMDB reviews [^21] https://ai.stanford.edu/~amaas/data/sentiment/, movies corpus [^22] https://zissou.infosci.cornell.edu/convokit/datasets/movie-corpus/, and Yelp 2013232323https://www.yelp.com/dataset/download. AI-generated content for these datasets was sourced from tweetHunter [^24] https://tweethunter.io/ and GPT, which was prompted to generate texts on specific topics such as American football, climate change, the Soccer World Cup, the Gaza conflict, U.S. politics, and vaccines.
The validation results are displayed in Table 7.
Platform Precision Recall
General 0.9064 0.5927
Facebook 0.8723 0.6686
Twitter 0.8935 0.5613
Telegram 0.9598 0.5956
Reddit 0.895 0.5368
Additional set 0.9697 0.5
Table 7. Precision and Recall for each platform
(a) Facebook.
(b) Twitter.
(c) Telegram.
Figure 5. Thresholds grid search for filtering similarity graph: x-axis edge similarity quantile, y-axis node centrality quantile. z-axis is the minimum graph density across all connected components of the filtered co-url similarity graph. The green square corresponds to the selected thresholds.
3. Grid Search for Similarity and Centrality Thresholds
To identify coordinated accounts in the co-URL similarity network, we filter edges and nodes based on cosine similarity and eigenvector centrality. Two nodes are connected if they co-share URLs, and the strength of these connections is represented by the cosine similarity between their TF-IDF vectors across the space of unique URLs.
We employ two filtering techniques commonly found in the literature: edge filtering, based on cosine similarity, and node filtering, based on eigenvector centrality. The thresholds for these filters are determined by the percentiles of the distributions of edge similarity and node centrality.
Assuming that coordinated accounts manifest as dense components in the similarity graph, we use the density of each connected component as a key indicator of coordination. To ensure robustness, we adopt a conservative approach, using the minimum density across all connected components after filtering as a comprehensive quality measure of the graph.
Figure 5 presents the results of our grid search across these parameters. The x-axis represents the percentile of edge similarity, while the y-axis corresponds to the percentile of eigenvector centrality. The z-axis indicates the minimum density of the similarity graph after filtering based on the selected x-y percentiles.
For each platform, we identified the optimal thresholds by observing sharp changes in minimum density and high overall density values. These threshold combinations define the coordinated accounts on each platform. Specifically:
  • For Facebook, we selected (50%, 45%) as the minimum density increased from 0.73 and 0.80 to 0.90.
  • For Twitter, we chose (85%, 99%), consistent with previous studies, leading to an increase from 0.74 to 0.99.
  • For Reddit, we selected (85%, 80%), where the minimum density shifted from 0.81 and 0.89 to 0.96.
  • For Telegram, we used (99%, 99%), focusing on a significant jump from 0.18 to 0.87.
Other threshold combinations were evaluated qualitatively, yielding similar results. More quantitative methods for threshold selection are left for future work.
Figure 6. Thresholds grid search for filtering the Cross-platform similarity graph: x-axis edge similarity quantile, y-axis node centrality quantile. z-axis is the minimum graph density across all connected components of the filtered co-url similarity graph. The green square corresponds to the selected thresholds.
4. Characterization of coordinated users
We report the bios of the top 5 coordinated accounts by degree, at the time of writing, and the top 5 messages by total engagement written by coordinated accounts in each platform.
Bio
CV Lies & 5G Dangers - Discussions
Scammers may try to impersonate Real World News Channel.
Don’t trust any messages you get from them.
They are not us. Block and report them immediately.
Covering News, Military information, across Wiltshire & the
Southwest areas and for connecting people together uk_flag
Credence - Breaking News
Truth. Faith. Freedom.
Table 8. Current bios of top-5 coordinated channels in Telegram sorted by degree.
Bio
Have integrity. Be Just. Precinct Chairman.
Christian. Pro-Life. Conservative Political Activist. #CruzCrew
NO DM’S. Beautiful disaster.
Self proclaimed arbiter of great ideas.
Here to annoy the dumb asses.
NO LIST!! #Imvotingforafelon #animallover
CHRISTIAN, ULTRA MAGA, PATRIOT, CONSERVATIVE,
2A, FJB, TRUMP WON, TRUMP2016, TRUMP2020, TRUMP2024,
ELON MUSK, YELLOWSTONE, DALLAS COWBOYS
just a opinionated old cowgirl from WY
Traditional work ethic, traditional values
studied spirituality under a minister study 15 years worked 30+years.
I write as I see them, past present and maybe future.
don’t expect you to agree but think
Table 9. Current bios of top-5 coordinated accounts in Twitter sorted by degree.
Bio
#GayRightNews Bringing You Gay(and other)
News with a Right Leaning View. #GaysGoRight
#GaysForTrump Support Our President #GaysGoRight
This is the OFFICIAL Page of the #GoRight Movement
(due to loss of the previous pages we were forced to create a new page)
This is not the time to abandon the GOP it’s time to #GoRight and
create the Great American Revival of the GOP and America.
#GoRight and Join Us on #TheQiew
When Gays Go Right We Make America Great
Political organization https://gorightnews.com/
Table 10. Current bios of top-5 coordinated pages in Facebook sorted by degree.
Bio
All of the articles and opinions shared here are not the opinions of HVU.
Use your own discernment. All of HVU’s links in one olace
https://linktr.ee/highvibesup Family - Love - Learn
All of the articles and opinions shared here are not the opinions of HVU.
Use your own discernment. All of HVU’s links in one olace
https://linktr.ee/highvibesup Family - Love - Learn
Karli Channel The Real Karli and X only KarluskaP is real
everyone else is being sued for using copyright avi artwork
Please post Tartaria architecture, flat earth
Table 11. Current bios of top-5 cross-platform coordinated accounts sorted by degree.
Top-5 Messages
"US Secretary of State Antony Blinken Walking the Congressional Office Building Halls Surrounded by Police ’War criminal’, ’genocide secretary’ Protestors reminding him of his complicity in crimes against humanity. RealWorldNewsChannel RealWorldNewsChat"
"Its sickening, disgusting money laundering MF’s. In a meantime the people of United States are suffering! Can’t afford basic necessities!!"
"I like how zoinists shout poor me when they get caught out for there evil crimes Normal Jews hate zoinist Jews People of the world don’t hate Jews People of the world hate zoinists Jews n zoinists in general n what they stand for They think there the chosen ones and that gives them the rite to kill n genecide inercent people Zoinisium is hated by all races of the world And has no place for peace in humanity "
"Ukrainians in Galway are being advised to vote for two Nigerians and a Labour candidate to best serve their interests. ’Under no circumstances vote for radicals - Irish Freedom Party, Independent Ireland, The Irish People and all others who have the slogans ’Ireland for the Irish’. #ForeignInterference "
"Former CNN Anchor Chris Cuomo Admits to Suffering from a COVID Vaccine Injury ICYMI: There’s been a major shift in the official narrative. Follow Vigilant_News "
Table 12. Top-5 messages by total engagement for coordinated channels in Telegram
Top-5 Posts
"“WE THE PEOPLE” don’t play by your dictator rules. Just answer the damn question, crybaby. White House correspondents fire back after Biden snaps at reporter for refusing to ’play by the rules’"
"If Biden wins there will no more arguing because the will not be anything left to argue about. He’s handing the keys over to illegals and the new world order. In the end, the joke will be on the democrats. The WHO/WEF hates them too and knows they’re imbeciles. Everyone knows"
"’Shameful’: GOP lawmaker shreds ’AWOL’ Biden for throwing Jews ’under the bus’ amid anti-Israel protests If this president, so-called president doesn’t personify evil destruction division of this country, I don’t know what or who would !"
"Biden DHS docs suggested Trump supporters, military and religious people are likely violent terror threats. HaHaHa! I guess we also have Santa Claus/Easter bunny over here in training. Give me a break, we are trying to save this country from the bad guys!"
"DeSantis spox dunks on NYT ’fact-check’ on terrorists entering southern border: ’Awaiting your correction’ They’ve been entering for last 3 years. Biden admin has no damn idea who is coming in. Even ones they think they know are using others’ identity"
Table 13. Top-5 posts by total engagement for coordinated accounts in Twitter
Top-2 Posts
"Youth Vote Shifts Toward Trump in 2024 Election https://gorightnews.com/youth-vote-shifts-toward-trump-in-2024-election/ Maybe the kids are alright… #GoRightNews Recent polls indicate a surprising trend: young voters are warming to Donald Trump in the 2024 presidential election. The question arises: Are these voters aligning with Trump’s policies, or is President Biden driving them away? The answer might lie in a combination of both factors. Polling Data Reveals Shift According to the latest New York Times poll, young voters aged 18 to 29 favor Biden by a slim margin of two points, 47% to 45%. A Quinnipiac poll shows Trump leading Biden among voters aged 18 to 34, 48% to 47%. This is a stark contrast to the 2020 presidential election, where Joe Biden secured the youth vote by a significant 24%. The last time a Republican won this demographic was in 1988. Biden’s Struggling Message Aidan Kohn-Murphy, founder of Gen Z for Change—a group that supported Biden in 2020—stated in the Washington Post, “Biden is out of step with young people on a number of key issues." Key issues where Biden seems to be losing support include: The War in Gaza: A majority of young voters, 51%, support the Palestinians, while only 15% support Israel. TikTok Ban: Biden’s support for a TikTok ban is perceived as an attack on free speech. Consequently, 67% of Gen Z voters say this makes them less likely to vote for him. Economic Challenges: High inflation and interest rates have made essential costs like food and housing unaffordable for young people entering the workforce or trying to purchase their first home. Trump’s Resonating Message The 45th president’s message appears to be resonating with young voters for several reasons: Gaza Conflict: While Trump supports Israel, he promises to bring a peaceful end to the conflict, citing his administration’s four years of peace as evidence of his capability. TikTok Engagement: Trump recently joined TikTok, emphasizing that Biden wants to shut the platform down, thus appealing to younger users. Economic Performance: When asked about Trump’s handling of the economy, 65% of young voters approved, compared to just 33% for Biden. Social Media Dynamics According to CredoIQ, a social media analytics firm, nearly 25% of the top left-leaning content creators on TikTok have posted anti-Biden content in the first four months of 2024, garnering over 100 million views. This content is often created by young, non-white liberals who share the belief that the U.S. Government, and specifically Joe Biden, aims to restrict free speech and information flow. Trump’s Adaptation to New Media In 2016, Trump broke political norms with an ’anyplace, anytime’ approach, dominating cable TV. For the 2024 campaign, he has adapted this strategy to new media, appearing on podcasts, YouTube shows, and attending live events such as UFC and Formula 1. He has also made campaign stops at local venues, including bodegas, firehouses, and even in the South Bronx. The shift in the youth vote suggests a significant realignment in political affiliations, one that underscores the importance of addressing the issues most pertinent to young Americans. This analysis highlights the potential impact of these changing dynamics on the upcoming presidential election. [Source: Washington Post, Axios] https://archive.is/kk3HU https://www.axios.com/2024/06/13/trump-election-young-voters-polling #GoRightNews Shared by Peter Boykin - American Political Commentator / Citizen Journalist / Activist / Constitutionalist for Liberty Web: https://PeterBoykin.com Kick: http://Kick.com/PeterBoykin YouTube: https://youtube.com/PeterBoykinForAmerica Twitter: https://twitter.com/GoRightNews Telegram: http://t.me/realpeterboykin Rumble: http://Rumble.com/GoRightNews Like the Content? Please Support! - Go Right News: Stripe: https://gorightnews.com/donations/support-gorightnews/ Cash App: http://cash.app/\$PeterBoykin1"
"Rising Costs Highlight Challenges for American Families Cost of rent, energy, and other essentials surged in May In an alarming trend, the cost of essentials such as rent, energy, and groceries continues to surge, underscoring the persistent financial challenges faced by American families. While the overall Consumer Price Index (CPI) showed a slight stabilization, the specifics reveal a stark reality of escalating living expenses in our Constitutional Republic. Analyzing the Numbers The CPI indicated a 3.3% rise in overall inflation compared to the previous year. Although this marks a slight decrease from April’s 3.4% and a significant drop from the 9.1% peak in June 2022, it remains well above the Federal Reserve’s target rate of 2%. This persistent inflation underscores the ongoing economic strain on American households. Historical Context It’s noteworthy that during the four years of Donald Trump’s presidency, the average inflation rate was maintained at a modest 1.9%. This comparison highlights a more stable economic period and suggests a need for policies that can effectively manage inflation without compromising the financial well-being of citizens. Essential Expenses on the Rise A closer examination of the May report reveals substantial increases in essential costs: Rent: Up by 5.4% Mortgage: Up by 5.6% Hospital services: Up by 7.2% Car insurance: Up by 20.3% Electricity: Up by 5.9% Ground beef: Up by 4.9% Steak: Up by 5.7% Bacon: Up by 6.9% Hot dogs: Up by 7.3% These increases in essential goods and services strain the budgets of American families, making everyday living increasingly unaffordable. Public Sentiment A recent poll reflects the public’s discontent, with only 31% of voters approving of President Biden’s handling of inflation, while a significant 61% disapprove. This sentiment underscores the urgent need for effective economic policies that address the real concerns of the populace. Administration’s Response The Biden administration continues to assert progress in combating inflation. A statement from the White House on social media claimed: "Today’s report shows continued progress in lowering inflation. President Biden knows that costs are still too high for many families and we still have a lot more to do. That’s why he will keep fighting to lower drug costs, grocery prices, and energy bills." Critical Perspective However, many Americans find these assurances lacking. The everyday experience at grocery stores and gas stations starkly contrasts with the administration’s optimistic declarations, leading to a disconnect between government rhetoric and public reality. As a Constitutional Republic dedicated to ensuring democracy and the well-being of its citizens, it is imperative that our government implements policies that stabilize the economy and reduce the financial burden on American families. The rising costs of essential goods and services are a pressing concern that requires immediate and effective action to safeguard the economic future of our nation. [Source: Poll, Whitehouse on X, BLS] https://d3nkl3psvxxpe9.cloudfront.net/documents/econTabReport_maqVHQt.pdf https://x.com/WhiteHouse/status/1800957041390792843 https://www.bls.gov/news.release/pdf/cpi.pdf #GoRightNews Shared by Peter Boykin - American Political Commentator / Citizen Journalist / Activist / Constitutionalist for Liberty Web: https://PeterBoykin.com Kick: http://Kick.com/PeterBoykin YouTube: https://youtube.com/PeterBoykinForAmerica Twitter: https://twitter.com/GoRightNews Telegram: http://t.me/realpeterboykin Rumble: http://Rumble.com/GoRightNews Like the Content? Please Support! - Go Right News: Stripe: https://gorightnews.com/donations/support-gorightnews/ Cash App: http://cash.app/\$PeterBoykin1"
Table 14. Top-2 posts by total engagement for coordinated pages in Facebook
Top-5 Messages
"!You are a Golden Child! A Golden Child 94 Unicorn 94 Third Eye 94 Covenant 94 Harmony 94 Praise God 94 Divine Gene 94 Nine Six 94 Blue Eyes 94 Indigo Child 94 Ultra Maga 94 John John 94 White Hat 94 American Eagle 94 Carry On 94 New Earth 94 Pineapple 94 PassionForFruit"
"#CAWildfireSituationUpdate as of 6-16-2024, 5:03 P.M. PST. #PostFire - 12,266 acres, 2% containment #HesperiaFire - 1,330, 7% #JunesFire - 1,076, 70% #JacksonFire - 876, ?#HernandezFire - 600, 25% #MaxFire - 500, 0% #LisaFire - 350, 0% #PointFire 150, 0%"
"What is it like to be wiser than your Creator? Job tried that and later repented in dust and ashes. Job was a wise and blessed man of faith."
"What was the true history behind all of these melted ruins? Certainly not the result of erosion, but perhaps a plasma storm? The results of a sudden flip in our polar magnetic electric field perhaps."
"WAVELAND, Mississippi A boil advisory was issued Mon afternoon dt burst in the main water line. NASHVILLE, Indiana A boil advisory was issued Mon after a water main break on Honeysuckle Ln. MONTICELLO, Kentucky A boil advisory was issued Mon for the East Hwy 92 area dt a water main break. MILAN, Ohio Seminary Rd bw Perrin Rd/Broad St in Milan is closed dt a water main break. SASKATOON, SK - Canada North Park Wilson School is closed dt a water main break. NEW CASTLE CO., Delaware Shipley Rd at Foulk Rd was closed Mon night dt a water main break. MONTGOMERY CO., Maryland Dt a water main break Mon, a portion of the NW Branch Stream Valley Park near Highwood Terrace was undergoing repairs. SIOUX FALLS, South Dakota A water main break Mon afternoon caused flooding in downtown. BOURBON CO., Kansas There is a water main break at the Bourbon Co Transfer Station. FORTUNA, California Water will be shut off Wed to repair a broken water main on S Fortuna Blvd."
Table 15. Top-5 posts by total engagement for cross-platform coordinated accounts
5. Topic Analysis
We report the most representative posts, tweets and messages—i.e., Facebook, Twitter/$\mathbb{X}$, Telegram—for each topic in Table 16, 17 and 18, as computed by BERTopic.
Topic Representative Posts
late-show
against-trump "Stephen Colbert Speaks Truth to Trump’s Most Damaging Lie - #stephencolbert #TheLateShow #Trump #DonaldTrump #MAGA #immigrants" "Stephen Colbert Speaks Truth to Trump’s Most Damaging Lie - Rick Strom #stephencolbert #TheLateShow #Trump #DonaldTrump #MAGA #immigrants"
barron-trump-rnc "Former President Donald Trump said his youngest son Barron Trump was 17-years-old in an interview, when in fact he is 18 and a nominate delegate for the RNC." "Former President Donald Trump’s 18-year-old son, Barron Trump, is serving as a delegate to the Republican National Convention."
2020-election-acceptance "Republican National Committee co-chair Lara Trump said it’s ’obvious’ that Donald Trump has accepted the 2020 election results, when in fact, he has not." "Republican National Committee co-chair Lara Trump said it’s ’obvious’ that Donald Trump has accepted the 2020 election results, when in fact, he has not."
trump-trial-dayoff "Judge Juan Merchan gave Donald Trump the day off from his hush money trial to attend Barron Trump’s graduation ceremony, but evidence suggests that he is attending political fundraiser." "Judge Juan Merchan gave Donald Trump the day off from his hush money trial to attend Barron Trump’s graduation ceremony, but evidence suggests that he is attending political fundraiser."
jacked-up-biden "Rep. Greg Murphy, a Republican, says he has evidence that President Joe Biden was using "something" before his popular State of the Union address." "Rep. Greg Murphy, a Republican, says he has evidence that President Joe Biden was using "something" before his popular State of the Union address."
trump-classified-documents "More bombshell evidence dropped against former President Donald Trump in his classified documents case, just as Judge Aileen Cannon delayed the trial indefinitely." "More bombshell evidence dropped against former President Donald Trump in his classified documents case, just as Judge Aileen Cannon delayed the trial indefinitely."
rfk "Exciting Announcement! Join us for a special interview with independent presidential candidate RFK Jr. on Monday, June 17, immediately following The Young Turks! Tune in on tyt.com/live YouTube, Facebook, or Twitch! TYT members get exclusive access to a members-only Bonus Episode after the interview, where Cenk will ask even more personal questions and dive deeper into the issues! Sign up as a member today at to join the discussion! Don’t miss this unique opportunity to engage with crucial political dialogue and watch TYT hold a presidential candidate accountable! See you there!" "Exciting Announcement! Join us for a special interview with independent presidential candidate RFK Jr. on Monday, June 17, immediately following The Young Turks! Tune in on tyt.com/live YouTube, Facebook, or Twitch! TYT members get exclusive access to a members-only Bonus Episode after the interview, where Cenk will ask even more personal questions and dive deeper into the issues! Sign up as a member today at to join the discussion! Don’t miss this unique opportunity to engage with crucial political dialogue and watch TYT hold a presidential candidate accountable! See you there!"
coulter-ramaswamy "Conservative Ann Coulter told former Republican presidential candidate Vivek Ramaswamy said while she agrees with him on many issues, she wouldn’t vote for him because he’s “an Indian.”" "Conservative Ann Coulter told former Republican presidential candidate Vivek Ramaswamy said while she agrees with him on many issues, she wouldn’t vote for him because he’s “an Indian.”"
trump-falsifying-records "Fox News’ Jeanine Pirro and conservative commentator Charlie Kirk are melting down over former President Donald Trump’s 34-count conviction for falsifying business records in his hush money trial."
bannon-sentence "Donald Trump adviser Steve Bannon proclaimed that "́no one" will shut him up as he was ordered to report to prison for refusing to testify before congress."
Table 16. Topics and their representative posts on Facebook
Topic Representative Tweets
rfk "If Kennedy iselected it will cost the evil doers over a trillion dollars in the first year. If Kennedy is in the debate it will show Biden as sinile, Trump as weak and Bobby will be elected. So they won’t let Bobby in but will compensate CNN and anyone that might go to jail." "I’ve never been more convinced that both Trump and Biden fear Kennedy. The uniparty has no soul, and they hate democracy. We will never forget their attempts to silence us. Kennedy FTW"
debate "Biden’s really not up to debate is what this means. Whenever a question is asked, Biden’s promoters will either tell the answer in an earplug or use a prompter screen on the podium so Biden can read the answer. This is funny." "Trump has guts though. They tried to make it as unfriendly for him as possible, and he still is going to do it. I’m sure they were hoping he would say no way, and then Biden could brag about Trump being afraid to debate him."
covid-vaccine "Remember this on election day, remember All the "suddenly died" family and friends! Biden did this, he made vaccines mandatory! Just one of many bad Biden decisions!" "Trump was during Pandemic; FAILED to protect the American people when he knew damn well how infectious; deadly Covid Virus was, instead Trump told America "Covid was a Hoax"; failed to create a National Suppression Plan to save lives 700KAmericans DIED under Trump!"
bowman-latimer "Bowman called Biden a liar. Bowman voted against the Infrastructure Act. Bowman even voted against the Debt Limit increase that could have jeopardized Social Security. Bowman did not earn another term as a Democrat." "Bowman is his worst enemy. His and Rape denying has turned his district against him. His Votes AGAINST the Biden Administration’s Progressive policies are NOT helping his district! Vote !"
border-security "with your lies. It is that BLOCKED Bipartisan Border Security Bill bc told GOP to "Kill The Bill". publicly stated it. The Border Crisis is NOW all on for Blocking Border Bill to SECURE THE BORDER!" "This bill was a sham and would not close border,but would allow more to take American jobs!! Donalds reiterates why GOP rejected ‘bipartisan’ border bill to head off potential debate talking point"
epstein-files "His flights with Epstein were with family present and years before he had an island. He’s also the only one saying anything about releasing the logs, Biden sure isn’t, and Trump took umbrage when DeSantis started advocating for their release for some strange reason…" " just another day in ’Murica, home of the child sex slaves. Trump made these folks feel they deserved Matt Gaetz’ job, Trumps’ job or Epstein’s job if they just ’networked’ enough. Investigate the real ’satanic’ sex cult that police and politicians hide."
student-loan "This administration and president Biden seem to be answerable to nobody between giving money to student loan forgiveness even after the court ruled against them not getting Kennedy secret service protectiin, the list goes on no accountability." ""Summary: We estimate that President Bidenś recently announced "New Plans" to provide relief to student borrowers will cost \$84 billion, in addition to the \$475 billion that we previously estimated for President Bidenś SAVE plan.""
hunter-biden "Hunter Biden’s laptop has nothing to do with the false Electoral College votes sent in for the 2020 election. Also, Alexander Smirnov (an FBI informant) was INDICTED for his false accusations about Barisma and the Bidens." "Actions speak louder than how they phrased something – especially when the actions are repeated over and over again. Their lies are proven in this week’s Hunter Biden trial where the laptop is used as evidence."
blacks-against-biden "People aren’t paying attention and don’t expect him to be the President and don’t know about Biden’s accomplishments. That’s what the polls show now but that’s gonna change. Trump isn’t gonna win 20% of the black vote. No were here near it" "’At this stage in his presidency (3.5 years in) Biden must stop trying to convince Black voters his policies have made the US a significantly better place. Instead he should spend everything he’s got highlighting how anti-Black Trump is and how quickly any past gains will be lost."
Table 17. Topics and their representative posts on Twitter
Topic Representative Messages
vaccine-skepticism "BREAKING: US doctors and scientists are currently investigating whether the COVID-19 virus is to blame for an “unusual pattern” of rare and deadly cancers that have been popping up in the wake of the pandemic.GeneralMCNews" "You, the unvaccinated are a special group of people. You stood your ground. You did not let the pressures of society sway you. You withstood the harshest discrimination seen in modern times. You were excluded from society. Some of you lost friendships and relationships with family members. You are the best that society has to offer. You used your analytical judgment and common sense. You are the tree that withstood the hurricane. You are superheroes! CovidVaccineTruth CovidVaccineTruthChat"
[…]
flat-earth-conspiracy "The Encyclopedia of Freemasonry reveals that the Earth is indeed flat. Organizations like The Freemasons know the truth about God and our flat earth but they choose to keep it hidden." "Flat Earth was the cosmology of our ancestors who had direct contact with the Divine because they were closer to the Divine than we are now. That is how the ancient’s knew things despite not having access to technology we have now."
[…]
Table 18. Topics and their representative posts on Telegram